Q2.Find the derivative of the following functions: 6 point a) y x5*x7 b) y 1/x-5 c) y = (-)8x6 d) y x.logx
Answer #1

Q y = q** at til =y= 95+7 = gold twi we have y zorle y differentiating (2) with respect to of, we have play an ghen? = 120 *(C) y = a (-) pour = y = -8070 (4) differentiating equation (4) with Lespect to a, we have po e (-8374) =-8d170 = 1-8)(-6), pda s** **(15) + Nang Ba 7 427) { by chain tule of differentication) = 34*0#594 + her, or +1492) = 504-5+4+4930* logget = 500+